# P.O. Box 15500, 00076 Aalto, Finland

### v1

@@ Abstract
A key issue with standard network measures of closeness and betweenness centrality is that they rely on the shortest paths between nodes within the network structure, whereas the degree centrality only reveals the immediate neighborhood of a node. Furthermore, many measures found in the literature do not accurately represent the physical or probabilistic characteristics of nodal centrality, network flow, and other salient properties. For example, recurrent spreading in a network is often overlooked by these metrics. Standard centrality measures have limitations, being optimal for one application but not for others. Here, we present new metrics based on our influence spreading model to characterize network structure for various network science applications. These probabilistic measures account for all feasible walks and cycles in the network. We compare our new metrics with the standard metrics in terms of the node rankings given by different centrality measures, by examining scatter plots, and by using the Pearson correlation coefficient. In the influence spreading model, we define the in-centrality measure to characterize how central a node is as a target of influence by other nodes and the out-centrality measure to characterize how central a node is as a source of influence on other nodes. Our results show that the influence spreading betweenness centrality reveals the importance of alternative routes while maintaining similarity to standard betweenness centrality.
@@ 1 Introduction
Network science is a rapidly expanding field of research that focuses on investigating the structure and dynamics of interdependent phenomena in complex systems of various scientific disciplines. This approach has been applied especially in studies of social science and communication networks, but also in studies of biological networks and its numerous subfields. In these fields of research, various modeling and other computational methodologies have become important in characterizing and analyzing the properties of these networks. The methods used in network science comprise a range of approaches, including statistical techniques, visualization methods, and various metrics that describe the structure, properties, and processes in networks. An important area of research focuses on clusters and communities in networks, where numerous methods for detecting communities have been proposed. In addition, basic tools, such as network centrality metrics, help measure the importance of individual nodes within the network structure.
In this article, we will frequently refer to paths, walks, and cycles. Based on [bender2010] and [bovet2022centralities], we define a walk on a network as an alternating sequence of nodes and edges, where every node is incident to both the edges that come before and after it in the sequence, and a path on a network is a walk where all nodes and edges are distinct. It is important to emphasize that all paths are walks, but all walks are not paths. A cycle is a walk where edges are required to be distinct, and the only repeated nodes are the first and the last node of the walk [bender2010].
Standard network centrality metrics, such as degree centrality, closeness centrality, and betweenness centrality, have become common in network science. The degree centrality $C_{D}(v)$ of node $v$ measures the number of neighbors $k_{v}$ that the node has with respect to the total possible number of neighbors [saxena2020centrality], as follows
@@
where $N$ is the total number of nodes in the network. The closeness centrality $C_{C}(v)$ of a node $v$ measures how close the node is to all other nodes in the network [saxena2020centrality]. It is calculated as the reciprocal of the average length of the shortest paths between node $v$ and all other nodes in the graph:
@@
The betweenness centrality $b(v)$ measures how frequently a node $v$ lies on the shortest path between two other nodes in the network:
@@
where $\sigma_{st}$ represents the total number of shortest paths from node $s$ to node $t$, while $\sigma_{st}(v)$ denotes the number of paths that pass through node $v$  [freeman1977set]. The high value of betweenness centrality indicates that a node often acts as a gatekeeper of information or resources. Numerous variants of both the closeness and betweenness centrality can be found in the literature [saxena2020centrality, bockholt2018, bockholt2021].
It should be noted that standard centrality measures have limitations, as they may be optimal for one application but suboptimal for another. In order to choose metrics, characteristics to consider are whether the network is directed or undirected and what is being analyzed, such as nodes with high degree (hubs) or those with mediating control, influential nodes, or the smallest set of nodes that can spread information [bovet2022centralities]. Moreover, it is important to assess variations in the ranking of different centrality measures to understand their contextual importance.
@@
In the present study and being inspired by the influence spreading model [kuikkaSciRep], we propose new metrics to describe consistent network characteristics, which are designed to be applicable in a wide range of network science scenarios. These new metrics consider all feasible paths, walks, and cycles within the network structure. In Section 2, we discuss how the influence spreading model is related to several important characteristics of networks. In Section 3, we give an overview of the application areas of the influence spreading model. In Section 4, we briefly discuss the research related to our study. In Section 5, we introduce the influence spreading model, give a motivating example of how the model works, and finally introduce our new metrics. In Section 6, we present the results of how the new metrics behave on multiple data sets compared to standard centrality metrics. In Section 7, we discuss our research results and how they relate to network modeling. Finally, in Section 8 we summarize our findings.
2 Network Structure and Dynamics
Here we highlight key characteristics of networks with new metrics that describe their structures, focusing on nodes and edges [kuikkaSciRep]. Additionally, we describe network flow as a spreading process of non-exclusive events within the network structure.
In the influence spreading model, weights are assigned to both nodes and edges. In this probabilistic framework, the weight of an edge represents the probability of influence or information passing through that edge, while the weight of a node indicates the probability of the node changing its state and enabling the passage of influence or information. These two properties can also be modeled as distinct features.
In the present study, the edges are considered to be directed, unidirectional, or bidirectional. For the bidirectional edges, it is possible to assign different weights for each direction. Undirected edges can be modeled by setting the same edge weight to each direction in a bidirectional edge. Then removing an edge or node from the network structure is equivalent to setting its weight to zero.
One of the most important features of our model is that cycles (or loops) [jiang2025identifying] are allowed in the network structure without restrictions. Moreover, these cycles can exist within other cycles and bidirectional links can be included in cycles, allowing flow to occur in both directions. An example of this is brain networks, where the cyclic propagation of signals can take place.
The definitions of standard network centrality measures, such as closeness centrality and betweenness centrality, count only geodesic or shortest paths between two nodes in the network. This is a limitation because it overlooks other paths and walks that can significantly influence various centrality properties and flow characteristics in the network. Our proposed centrality metrics address this issue by considering all possible walks that are shorter than a maximum walk length, which is defined as a model parameter. When the edge probabilities are low, due to decay effects, the model accounts for all relevant walks with the accuracy of the computations.
Centrality measures inherently imply specific patterns of flow in networks [borgatti2005]. This is true even for degree centrality, unless the network graph is used solely for visualization purposes. Two main extreme types of processes occur in networks: spreading along self-avoiding walks and spreading along node-state-independent walks. In the case of self-avoiding walks, each node can only be visited once, whereas in the case of node-state-independent walks, there are no such restrictions. Evidently, cycles cannot appear on self-avoiding walks.
Network centrality metrics and community detection methods have been developed both for local and global network structures  [qiu2021ranking]. Here the term global refers to approaches that take into account the entire network. In many cases, global methods address the limitations found in standard metrics or other measures proposed in the literature. Many of them are based on existing techniques, enhance them with improvements or modifications, or use combinations of standard metrics [qiu2021ranking, meng2025development, kouam2024novel].
3 Applications
In this section, we highlight some representative applications where our proposed metrics play a significant role. This list is not exhaustive, as network centrality metrics are crucial in most network analysis studies.
Social network analysis can help us find influential leaders in a social group, trace the source of new ideas and innovations, and assess the speed of information spreading [brass2022]. In social network analysis, an important application is the detection of the source or sources of rumor, misinformation, and disinformation [shelke2019]. The characteristics of influence spreading involve complex social contagion mechanisms, where multiple interactions are necessary for an opinion or behavior change to occur [centola2007]. In contrast, information or persuasive comments may require just a single interaction to propagate the news. It is important to note that we use the term complex social contagion to refer to Damon Centola’s concept of complex contagion [centola2007], where multiple influences are needed for an individual to change their opinion or behavior. In section 5.1, we describe in detail a different meaning for the term complex contagion in the context of the influence spreading model.
Modeling the spread of epidemics has become a significant area of research due to the COVID-19 pandemic. New approaches have been developed to describe both the spread and containment of the virus. Detailed modeling at the organizational level has allowed for targeted isolation of specific units within organizations and has enabled the investigation of how individuals in different contact networks contribute to spreading the virus [kuikkaSciRep]. Commonly used metrics to summarize model outcomes include network centrality, betweenness, and the basic reproduction number.
Communication networks are usually evaluated using various metrics that indicate their connectivity and reliability. In this context, circular propagation is not relevant, which allows for the use of self-avoiding walk models. When investigating cyber threats and attacks, it is important to model at both the enterprise and cross-enterprise levels. The use of various technical attack techniques and multi-layer approaches highlights the necessity for detailed modeling of these environments [kuikka2025network].
@@ 4 Related Work
Centrality measures based on shortest paths make simplistic assumptions of how information, traffic or other network processes flow through a network [freeman1991, borgatti2005, bockholt2021]. The spread of information can either randomly or by intention take a longer and indirect route with many intermediaries [freeman1991]. Stephen Borgatti noted how a flow process can use unrestricted walks and where there can either be transfer (no duplication), serial duplication or parallel duplication depending on what the flow process consists of [borgatti2005]. Bockholt and Zweig noted that the use of certain centrality measures contains an implicit assumption on how the network process flows through the network [bockholt2018]. In another publication, the authors also note that standard betweenness centrality is not the best at identifying influential nodes in an information spreading process [bockholt2021]. However, Freeman [freeman1978] and Krackhardt [krackhardt1990] argued that standard betweenness centrality is closely associated with social power, since a node with high betweenness centrality can control communication and the flow of information. Krackhardt also noted that a high standard betweenness centrality value can indicate that a node with high betweenness has access to a variety of non-redundant sources of information.
The relationship between degree centrality and closeness centrality was recently studied in [evans2022]. The authors found a non-linear relationship between degree centrality and closeness centrality. Degree, closeness, betweenness, eigenvector, and Katz centralities were mathematically examined with respect to whether these measures fulfill the authors’ three basic requirements for centrality measures in [landherr2010]. The correlation between community-aware centrality measures and classical centrality measures was studied in [rajeh2021].
Freeman discusses the notion that the center of a star-shaped network or the hub of a wheel-shaped network is structurally more central than any other node in any network of similar size [freeman1978]. However, this is more of a convincing argument than an exact mathematical definition. A comprehensive survey of centrality measures in complex networks gives a nice summary and overview of centrality measures [saxena2020centrality]. The survey highlights how degree centrality is a centrality measure based on local information, whereas closeness centrality and betweenness centrality make use of global knowledge of the network. Bovet and Makse give an overview of centrality measures based on whether they are based on shortest paths, walks, random walks, or non-backtracking walks [bovet2022centralities]. They note that degree centrality can actually be understood as the number of walks of length one starting from a node.
In a foundational work, Friedkin proposed a process model of social influence from which three centrality measures were defined [friedkin1991]. More recently, many new centrality measures have been proposed to identify influential nodes in complex networks [esfandiari2024collaborative, sadhu2024normalized, ghorbani2024opinion, bendahman2024unveiling, chiranjeevi2023icdc, qiu2021ranking].
@@ 5.1 Influence Spreading Model
Our influence spreading model, previously published in [kuikkaSciRep], provides a detailed description of network structure and non-conserving probabilistic influence flow within that structure. The edges in the network can be directed or bidirectional, and both nodes and edges are assigned individual weights or probabilities for transmitting influence. The model supports both circular and recurrent spreading throughout the network. Additionally, a version of the model has been developed to implement only self-avoiding walks, meaning that each node can appear only once on a given walk. Furthermore, the model can account for breakthrough effects, enabling us to describe scenarios that lie between these two extreme cases.
The influence spreading model allows for two distinct types of spreading, namely complex contagion and simple contagion. Complex contagion allows unrestricted self-intersecting walks and cycles, with the only limitation for walks being a specified maximum walk length, and the target node cannot self-intersect. However, simple contagion does not allow for self-intersecting walks or cycles. In complex contagion, the focus is on the first arrival or the first influence on a target node. Before reaching the target node, the walks are limited by the specified maximum length. The calculation for each walk ends when it reaches the target node. In this article, we focus only on complex contagion since we wish to study centrality measures related to unrestricted paths, walks, and cycles and compare them to standard centrality measures.
@@
The complete logic of calculating influence spreading probabilities was originally presented in [kuikka2022efficiency]. Depending on the context, the original article often used the term path interchangeably to mean either a path or a walk as defined in Section 1. In this article, we will aim to use more exact terminology that aligns with existing graph theory terminology.
The full influence spreading model specifies a time-dependent component in addition to edge and node probabilities [kuikka2022efficiency]. In the full model, this time-dependent probability component approaches one as time grows unbounded. In this article, we assume that the time component probability equals one. By setting the time component probability to one, we can have a clearer focus on how the influence spreading model’s centrality measures behave as the edge probability is varied.
In [kuikka2022efficiency], a detailed algorithmic description was given on how different walks and their probabilities are combined. Next, we will explain the logic of the walk combination process and the algorithm. Understanding how the influence spreading probability is calculated when walks are only restricted by the maximum allowed length of the walk is crucial to understanding the results that are presented in Section 6.
The probabilities of two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are combined by using their longest common prefix (LCP) walk $\mathcal{L}_{3}$ and the rules of probability theory. The reasoning behind this is to ensure that the results can be interpreted as valid probabilities. If $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ do not share any common walk, then we set $\mathbb{P}(\mathcal{L}_{3})=1$, where we denote probability with $\mathbb{P}$. The influence spreading model assumes that two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are conditionally independent given a common prefix walk $\mathcal{L}_{3}$ i.e. $\mathbb{P}(\mathcal{L}_{1}\cap\mathcal{L}_{2}\mid\mathcal{L}_{3})=\mathbb{P}(\mathcal{L}_{1}\mid\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{2}\mid\mathcal{L}_{3})$. The formula for combining the probabilities of walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ given a common prefix walk $\mathcal{L}_{3}$ is derived as follows. We have
@@
where the second equality follows from walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ both sharing the common walk $\mathcal{L}_{3}$, the third equality follows from the definition of conditional probability, the fourth equality follows from assumed conditional independence, and the fifth equality follows from $\mathbb{P}(\mathcal{L}_{i})=\mathbb{P}(\mathcal{L}_{i}\cap\mathcal{L}_{3})=\mathbb{P}(\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{i}\mid\mathcal{L}_{3})$ for $i=1,2$.
Equation (4) for combining two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ forms the basis of the influence spreading algorithm 1. Algorithm 1 and a more efficient algorithm for calculating influence spreading were originally presented in [kuikka2022efficiency]. For ease of understanding, we present the version which explicitly lists all possible allowed walks, although it is computationally less efficient. The computational limitations of algorithm 1 were discussed at length in [kuikka2022efficiency]. In practice, we have used the more efficient algorithm in our experiments. It is important to emphasize that both algorithm versions produce the exact same results.

### v2

@@ Abstract
Traditional measures of closeness and betweenness centrality in networks rely on the shortest paths between nodes. Many standard metrics fail to accurately reflect the physical or probabilistic characteristics of nodal centrality and network flow, often overlooking processes such as cyclic and recurrent spreading. Here, we present new metrics based on our influence spreading model. These probabilistic measures consider all feasible paths, walks, and cycles within the network. We define in-centrality to assess how central a node is as a target of influence, and out-centrality for its role as a source of influence. We compare our metrics with standard ones by analyzing node rankings, using scatter plots, and calculating the Pearson correlation and Spearman’s rank correlation coefficients. Our findings show that the betweenness centrality defined by the influence spreading model emphasizes the importance of alternative routes while maintaining similarity to standard betweenness centrality.
@@ 1 Introduction
Network science is a rapidly growing field that investigates the structure and dynamics of interdependent phenomena in complex systems across various disciplines. It has been particularly useful in social sciences, communication networks, and biological networks. Key methodologies in this field include statistical analyses, visualization techniques, and metrics that describe network properties. A significant area of research is the identification of clusters and communities in networks, for which many detection methods have been developed.
In the analysis of networks, many methods focus on the structural properties of paths, walks, and cycles. A walk is an alternating sequence of nodes and edges, with each node connected to its preceding and following edges. A path is a specific type of walk in which all nodes and edges are distinct. Therefore, while all paths are walks, not all walks are paths. A cycle, on the other hand, is a walk in which the edges are distinct and only the first and last nodes may be repeated.
To characterize the local and global network structures, various centrality metrics have been developed [26]. Standard network centrality metrics, such as degree centrality, closeness centrality, and betweenness centrality, have become common in network science. The degree centrality $C_{D}(v)$ of the node $v$ measures the number of neighbors $k_{v}$ that the node has with respect to the total possible number of neighbors [28], reading
@@
where $N$ is the total number of nodes in the network. The closeness centrality $C_{C}(v)$ of a node $v$ measures how close the node is to all other nodes in the network [28]. It is calculated as the reciprocal of the average length of the shortest paths between node $v$ and all other network nodes
@@
The betweenness centrality $b(v)$ measures how frequently a node $v$ lies on the shortest path between two other nodes in the network
@@
where $\sigma_{st}$ represents the total number of shortest paths from node $s$ to node $t$, while $\sigma_{st}(v)$ denotes the number of paths that pass through node $v$. A high value of betweenness centrality indicates that a node acts often as a gatekeeper of information or resources. There are numerous variants of closeness and betweenness centrality measures that can be found in the literature [28, 13].
Standard centrality measures have limitations, being optimal for some applications but not for others [24, 3, 4]. When selecting metrics, it is crucial to consider whether the network is directed or undirected and what aspects are under analysis, such as hubs or influential nodes. Standard network centrality measures, like closeness and betweenness centrality, focus on the shortest paths between nodes, which limits their effectiveness by ignoring other significant paths and walks. Our proposed metrics consider all possible walks up to a specified maximum length, capturing relevant paths and walks even when edge probabilities are low due to decay effects.
Centrality measures imply specific flow patterns in networks [6]. Two main types of process occur in networks: self-avoiding walks, where nodes are visited only once, and node-state-independent walks, with no such restrictions. Social network analysis is used to identify influential leaders and trace the spread of ideas and innovations [8]. It also helps to assess information dissemination processes within networks. Influence may require multiple interactions to produce opinion changes, whereas information can spread through a single interaction [29], highlighting the need to account for different spreading mechanisms.
Modeling epidemic spread has gained importance since the COVID-19 pandemic, leading to new approaches to understanding spread and containment at the organizational level [22]. Commonly used outcome metrics include network centrality and reproduction numbers. Communication networks are assessed using metrics that indicate connectivity and reliability, with self-avoiding walk models being relevant, especially in cyber threat investigations. Detailed modeling at the enterprise and cross-enterprise levels is essential, given the complexity of attack techniques and environments [23].
@@
In the present study, we propose new metrics to describe consistent network characteristics, which are designed to be applicable for a wide range of network science scenarios [22]. These new metrics consider all feasible paths, walks, and cycles within the network structure. In Section 2, we briefly discuss the research related to our study. In Section 3, we introduce the influence spreading model, give a motivating example of how the model works, and finally introduce our new metrics. In Section 4, we present the results of how the new metrics behave on multiple data sets compared to standard centrality metrics. In Section 5, we discuss our research results and how they relate to network modeling. Finally, in Section 6 we summarize our findings.
@@ 2 Related Work
Centrality measures based on shortest paths make simplistic assumptions of how information, traffic or other network processes flow through a network [13, 6, 4]. The spread of information can occur either randomly or intentionally take a longer and indirect route with many intermediaries [13]. Stephen Borgatti noted that a flow process can utilize unrestricted walks, and the outcome can be either transfer (no duplication), serial duplication, or parallel duplication, depending on the nature of the flow process [6]. Bockholt and Zweig noted that the use of certain centrality measures contains an implicit assumption on how the network process flows through the network [3]. In another publication, the authors also note that standard betweenness centrality is not the best at identifying influential nodes in an information spreading process [4]. However, Freeman [13] and Krackhardt [20] argued that standard betweenness centrality is closely associated with social power, since a node with high betweenness centrality can control communication and the flow of information. Krackhardt also noted that a high standard betweenness centrality value can indicate that a node with high betweenness has access to a variety of non-redundant sources of information.
The relationship between degree centrality and closeness centrality was recently studied in [12]. The authors found a non-linear relationship between degree centrality and closeness centrality. Degree, closeness, betweenness, eigenvector, and Katz centralities were mathematically examined with respect to whether these measures fulfill the authors’ three basic requirements for centrality measures in [24]. The correlation between community-aware centrality measures and classical centrality measures was studied in [27].
A comprehensive survey of centrality measures in complex networks provides a nice summary and overview of centrality measures [28]. The survey highlights how degree centrality is a centrality measure based on local information, whereas closeness centrality and betweenness centrality make use of global knowledge of the network. Bovet and Makse give an overview of centrality measures based on whether they are based on shortest paths, walks or random walks [7]. They note that degree centrality can actually be understood as the number of walks of length one starting from a node.
@@ 3.1 Influence Spreading Model
Our influence spreading model, previously published in [22], provides a detailed description of the network structure and non-conserving probabilistic influence flow in that structure. The edges in the network can be directed or bidirectional, and both nodes and edges are assigned individual weights or probabilities to transmit influence. The model supports both circular and recurrent spreading throughout the network. In addition, a version of the model has been developed to implement only self-avoiding walks, meaning that each node can appear only once on a given walk. As the model can account for breakthrough effects, it enables us to describe scenarios that lie between these two extreme cases.
The influence spreading model allows for two distinct types of spreading, namely complex contagion and simple contagion. Complex contagion allows unrestricted self-intersecting walks and cycles, with the only limitations for walks being a specified maximum walk length and that the target node cannot self-intersect. However, simple contagion does not allow for self-intersecting walks or cycles. In complex contagion, the focus is on the first arrival or the first influence on a target node. Before reaching the target node, the walks are limited by the specified maximum length. The calculation for each walk ends when it reaches the target node. In this article, we focus only on complex contagion since we wish to study centrality measures related to unrestricted paths, walks, and cycles and compare them to standard centrality measures.
@@
The complete logic for calculating the probability of spreading the influence was originally presented in [21]. Depending on the context, the original article often used the term path interchangeably to mean either a path or a walk as defined in Section 1. In this article, we will aim to use more exact terminology that aligns with existing graph theory terminology.
The full influence spreading model specifies a time-dependent component in addition to edge and node probabilities [21]. In the full model, this time-dependent probability component approaches one as time grows unbounded. In this article, we assume that the probability of the time component is equal to one. By setting the time component probability to one, we can have a clearer focus on how the influence spreading model’s centrality measures behave as the edge probability is varied.
In [21], a detailed algorithmic description was given on how different walks and their probabilities are combined. Next, we will explain the logic of the combination process of walks and the algorithm. Understanding how the influence spreading probability is calculated when walks are only restricted by the maximum allowed length of the walk is crucial to understanding the results that are presented in Section 4.
The probabilities of two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are combined by using their longest common prefix (LCP) walk $\mathcal{L}_{3}$ and the rules of probability theory. The reason behind this is to ensure that the results can be interpreted as valid probabilities. If $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ do not share any common walk, then we set $\mathbb{P}(\mathcal{L}_{3})=1$, denoting the probability with $\mathbb{P}$. The influence spreading model assumes that two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are conditionally independent given a common prefix walk $\mathcal{L}_{3}$ i.e. $\mathbb{P}(\mathcal{L}_{1}\cap\mathcal{L}_{2}\mid\mathcal{L}_{3})=\mathbb{P}(\mathcal{L}_{1}\mid\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{2}\mid\mathcal{L}_{3})$. The formula for combining the probabilities of the walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ given a common prefix walk $\mathcal{L}_{3}$ is derived as follows
@@
where the second equality follows from the walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ both sharing the common walk $\mathcal{L}_{3}$, the third equality follows from the definition of conditional probability, the fourth equality follows from assumed conditional independence, and the fifth equality follows from $\mathbb{P}(\mathcal{L}_{i})=\mathbb{P}(\mathcal{L}_{i}\cap\mathcal{L}_{3})=\mathbb{P}(\mathcal{L}_{3})\mathbb{P}(\mathcal{L}_{i}\mid\mathcal{L}_{3})$ for $i=1,2$.
Equation (4) for combining two walks $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ forms the basis of the influence spreading algorithm 1. The algorithm 1 and a more efficient algorithm for calculating influence spreading were originally presented in [21]. For ease of understanding, we present the version that explicitly lists all possible allowed walks, although it is computationally less efficient. The computational limitations of algorithm 1 were discussed at length in [21]. In practice, we have used the more efficient algorithm in our experiments. It is important to emphasize that both algorithm versions produce the exact same results.